OpenStudy (anonymous):

w(cube)=27,w is rational or irrational

OpenStudy (anonymous):

\[\sqrt[3]{27}\]=3

OpenStudy (anonymous):

\[w^3=27\]Thus w=3, rational.

OpenStudy (anonymous):

rational

OpenStudy (anonymous):

\[\sqrt[3]{3x-2}=4,find the valueofx\]

OpenStudy (anonymous):

3x-2=4^3 3x-2=64 3x=66 x=22